OpenStudy (anonymous):

y=(x-2)^2 I need to put this in format -b/2a Can anyone help me with this?

OpenStudy (anonymous):

expand and complete the square...

OpenStudy (tkhunny):

Please don't expand and complete the square - unless you are really fond of driving around in circles. IF you were to multiply this out, you would determine that you already have what you seek. x = -b/2a is the axis of symmetry. The axis of symmetry is x = 2, read directly from the present form.

OpenStudy (anonymous):

So if I plug in 2 where x is I get 0

OpenStudy (tkhunny):

Okay. You have found the vertex. The vertex lives on the axis of symmetry. You have not actually stated what the problem is. Are we just playing with parabolas?

OpenStudy (anonymous):

yes parabolas so If i now place zero for x I can obtain a vertex of 0,0 correct?

OpenStudy (tkhunny):

Where did the 2 go? We already established that x = 2 is the axis of symmetry and (2,0) is the vertex. You still haven't suggested what it is you are supposed to be doing.

OpenStudy (anonymous):

Ok yes vertex is 2,0, now I need to obtain some random points to find out if the parabola is going up or down, but my dilemma is that I am still not in the format of -b/2a so I can plug in random numbers.

OpenStudy (anonymous):

nevermind I think I just figured it out

OpenStudy (tkhunny):

The form is of no consequence. Substitute any value you like in any form that it is.

OpenStudy (anonymous):

Thank you tkhunny

OpenStudy (tkhunny):

If you REALLY want to multiply it out, go ahead. y = (x-2)^2 = x^2 - 4x + 4